OpenStudy (anonymous):

solve by factoring: x^2 + 10x + 25 = 9

jimthompson5910 (jim_thompson5910):

x^2 + 10x + 25 = 9 x^2 + 10x + 25 - 9 = 0 x^2 + 10x + 16 = 0 (x+8)(x+2) = 0 x+8 = 0 or x+2 = 0 x = -8 or x = -2